the length of a rectangle is 5 units longer than its width. The perimeter of this rectangle is 74cm. determine the length and the width of this rectangle.
Answer by oberobic(2304)

L = length
W = width
P = perimeter = 2L + 2W = 2(L + W)
.
L = W+5
P = 74
.
2L + 2W = 74
substituting
2(W+5) + 2W = 74
2W + 10 + 2W = 74
4W = 64
W = 16
.
L = W+5 = 16+5 = 21
.
Checking
P = 74 ??
2(21) + 2(16) = 42 + 32 = 74
Yes.
.
Answer:
L = 21
W = 16
.
Done